OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point. y = 3x + 7; (2, 10)

OpenStudy (anonymous):

You know the slope, and you have a point

OpenStudy (anonymous):

Do you know point-slope form?

OpenStudy (anonymous):

\[y-_{y1}=m(x-_{x1})\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

So all you have to do is plug in the y coordinate for y1 and the x coordinate for x1, and you already have the slope, so plug that in for m

OpenStudy (anonymous):

um? y - 10 = 7( x - 2)

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

Did your teacher teach you point slope form yet?

OpenStudy (anonymous):

yes, but that wasnt a answer choice

OpenStudy (anonymous):

what are your choices? All i need to know is the type of form (standard, slope intercept)?

OpenStudy (anonymous):

A) y = 3x + 11 B) y = 3x + 4 C) y = -3x + 4 D) y = (-1/3)x + 4

OpenStudy (anonymous):

okay so it's slope intercept...

OpenStudy (anonymous):

What you can do is simplify point slope

OpenStudy (anonymous):

Im not understanding. Which one is the answer?

OpenStudy (anonymous):

y - 10 = 7( x - 2) What would you do first?

OpenStudy (anonymous):

to simplify the equation

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

All you need to do is simplify the point-slope, so distribute the 7

OpenStudy (anonymous):

Im confused...

OpenStudy (anonymous):

you should get y-10=7x-14

OpenStudy (mrnood):

your equation is not correc the slope is 3 not 7 go backt to y− y1 =m(x− x1 )

OpenStudy (anonymous):

oh i didnt notice that thanks mr nood xD

OpenStudy (anonymous):

He said it was correct?

OpenStudy (anonymous):

The point slope form equation is: y - 10 = 3( x - 2)

OpenStudy (anonymous):

so to turn it into slope intercept first distribute the 3

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

you should get y - 10 = 3x - 6)

OpenStudy (mrnood):

let him do the work - he will learn more that way

OpenStudy (anonymous):

thats not a answer choice though.

OpenStudy (anonymous):

That is only one of the steps

OpenStudy (anonymous):

Im a girl...

OpenStudy (anonymous):

Do you know how to simplify equations?

OpenStudy (mrnood):

it is one of th echoices - you just need to recognise it add 10 to each side...

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

Is is y = 3x + 4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@MrNood ?

OpenStudy (anonymous):

All you did was isolate y

OpenStudy (anonymous):

that is slope intercept form

OpenStudy (anonymous):

@MrNood left.

OpenStudy (anonymous):

um okay... hope its correct. Thanks

OpenStudy (anonymous):

It is correct, you can trust me.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

its B